please help me solve this equation: (-6/y)^3
Answer by Theo(3458)

(-6/y)^3 = (-6)^3/y^3 = -126/y^3

this is because (a/b)^3 = a^3/b^3

let a = 5 and let b = 9

(a/b)^3 = (5/9)^3 = .171467764 (use your calculator to solve)

(a/b)^3 = a^3/b^3 = 5^3/9^3 = 125/729 = .171467764

they are equivalent.

what you were asking me to solve is not an equation, but an expression.

(-6/y)^3 is an expression.

when you make it equal to something, then the whole thing becomes an equation.

example:

(-6/y)^3 = 99635

now you have an equation.

the expression on the left side of the equal sign equals the expression on the right side of the equal sign.